John and Mark shared $170. John spent 1/5 of his money and Mark spent $10 more than John. The amount of money John had left was twice as much as the amount Mark has left. How much money did John spend?
Answer by ankor@dixie-net.com(22740)

John and Mark shared $170.
j = John's original amt, m = Mark's
j + m = 170
or
m = 170-j
:
John spent 1/5 of his money and Mark spent $10 more than John.
John spent .2j and Mark spent (.2j+10)
:
The amount of money John had left was twice as much as the amount Mark has left.
j - .2j = 2(m - (.2j+10))
replace m with (170-j)
j -.2j = 2((170-j)-(.2j+10))
.8j = 2(170 - j - .2j - 10)
.8j = 2(-1.2j + 160)
.8j = -2.4j + 320
.8j + 2.4j = 320
3.2j = 320
j = 320/3.2
j = 100
How much money did John spend? .2(100) = $20
:
:
Check this out
j had 100 then m had 70
j spent 20 leaving him 80
m spent 30 leaving him 40
j left with twice as much at m